# Beyond the Request: Harnessing HTTP Response Headers for Cross-Browser Web Tracker Classification in an Imbalanced Setting

### v1

@@ 6.1. Intra-browser Performance
Table 4 shows the performance of our $\text{Chrome}_{\textbf{22}}$-trained classifiers and the baselines. In terms of Accuracy, Ensemble Trees (ET) lead with a score of 0.964, closely followed by Random Forest (RF) and Decision Trees (DT) with scores of 0.960 and 0.956, respectively.
The Log-Loss metric, which penalizes classifiers for incorrect predictions, shows that ET, RF, and DT also perform well, with ET having the lowest log loss, indicating a high level of confidence in its predictions. On the other hand, Gaussian Naive Bayes (GNB) exhibits a significantly higher log loss, which suggests that it is less confident in its predictions compared to the tree-based methods. The ROC-AUC metric further strengthens the position of ET, RF, and DT, as they all score above 0.94, indicating that they are good at differentiating between the positive and negative classes. The high MCC scores corroborate the findings from the other metrics, placing these three classifiers as the most robust performers across the board. In contrast, baseline models like LR are comparatively lower performance across these metrics.
The gradient boosting models, particularly XGBoost, have fairly high scores as reflected in an AUC-ROC score of 0.915 and an F1-Score of 0.875. They are able to handle the class imbalance due to the fairly high MCC and AUPRC scores. However, their precision and recall scores indicate room for improvement in identifying instances of the minority class.
Comparing the best performing classifiers ET and RF to the baseline models, shows that they perform significantly higher, except for the DT and its Log-Loss score. The t.ex-Graph classifiers display lower performance across all metrics.
@@ 6.2. Cross-browser and Longitudinal Performance
Figure 5 presents the cross-browser performance and longitudinal performance of our $\text{Chrome}_{\textbf{22}}$-trained classifiers. We focus on three selected classifiers, one baseline model, one ensemble tree-based model, and one ensemble tree-based model with gradient boosting.
@@
For Logistic Regression and Random Forest, $\text{Chrome}_{\textbf{22}}$ shows high performance across most metrics, which is expected as the models would be most adapted to the data they were trained on. The performance on $\text{Chrome}_{\textbf{23}}$, $\text{Firefox}_{\textbf{22}}$, and $\text{Brave}_{\textbf{22}}$ varies, with a general trend of decreasing performance, indicating potential overfitting to the training data or a lack of generalizability to other datasets.
The XGBoost model shows a similar trend, with high performance on the training set and slightly varied results on the test sets. However, the variations in the XGBoost model’s performance are less pronounced than in the Logistic Regression and Random Forest models, suggesting that XGBoost may be more robust or generalizable across different browser datasets.
@@
Overall, the figure indicates that while all models perform well on the dataset they were trained on, their ability to generalize to other data is varied. XGBoost appears to have the most consistent performance across different browser datasets, indicating a potential advantage in terms of generalization. The figure underscores the importance of cross-browser testing to ensure that models are not only tuned to a specific dataset but also maintain their performance across diverse scenarios.
@@
The Brave results show a high accuracy metric for the overall performance, but the F1-Score and MCC are low and represent a more realistic performance of the classifiers, which is expected due to the more considerable differences between the Chrome and Brave data.
@@
The longitudinal performance analysis, as depicted in Figure 5, indicates a slight decline in the efficacy of the classification models when transitioning from the $\text{Chrome}_{\textbf{22}}$ to the $\text{Chrome}_{\textbf{23}}$ dataset. This is evidenced by the contraction of the metric lines from $\text{Chrome}_{\textbf{22}}$ to $\text{Chrome}_{\textbf{23}}$ across the three models, suggesting a decrease in metrics such as AUC-ROC, Accuracy, and others. Notably, the XGBoost model demonstrates the smallest degree of performance reduction, underscoring its robustness and adaptability to changes over time. In contrast, both the Logistic Regression and Random Forest models exhibit a more pronounced drop, highlighting potential challenges in model generalization across browser version updates. This suggests that while all models are affected by evolving data characteristics, ensemble methods, particularly gradient boosting as implemented in XGBoost, may offer a more stable predictive performance in longitudinal settings. Applying isotonic calibration had only minor improvements compared to the uncalibrated models as shown in Figure 6.
@@ 7. Discussion
Tracking affects users across all three browsers. However, this varies to a certain degree, e.g. Brave had the least amount of tracking-related HTTP responses compared to Chrome and Firefox.
Nevertheless, an imbalance can be observed in all data sets, which might be a problem for specific ML models. Even though many different methods were developed to balance data sets, due to the moderate imbalance in Chrome and Firefox and the good performance, a resampling was not conducted to preserve the observed distribution. Still, the large number of observations provides sufficient HTTP response messages for each class label. Furthermore, several evaluation metrics were calculated and provided that are aimed at imbalanced data sets, e.g. BACC.
The classifiers performed similarly in Chrome and Firefox but were poor with Brave data. First and foremost, the application of binary-encoded features resulted in surprisingly good results, i.e. the mere presence of a header value is enough and extensive feature engineering is not required from a pure tracking detection standpoint. On the other hand, explaining the models is limited, except for model parameters, training time, or the calculated metrics. In addition, this method simplifies the models to a high degree, although HTTP header values can reveal more information about trackers and their differences from non-trackers. Nevertheless, the binary approach helped preserve the core information, namely whether a header is related to tracking.
The NB classifier did not perform as well as others, but that might be due to its limited applicability to the data sets. As mentioned above, NB is based on Bayes Theorem and assumes that features within a class are not correlated to each other. It can be assumed that a set of headers and thus features are correlated to a certain degree. These were not removed which is a violation of Bayes Theorem in a statistical sense could be argued, and thus, its applicability. On the contrary, a paper by Berrar (Berrar, 2018) argues that NB can perform well, even under these circumstances. Other implementations, e.g., the Categorical Naïve Bayes might have been a good fit for the training sets with categorical features (nb, 2023).
@@
Filter lists are reliable and effective when it comes to blocking trackers (Mazel et al., 2019), and are one of the oldest methods, e.g. EasyList was created in 2005 (eas, 2023a). They present a popular ground truth to label data sets for training ML models. However, they have to be constantly updated, which commonly happens manually. A longitudinal study of nine years by Alrizah et al. (Alrizah et al., 2019) showed that human errors, e.g. acting in a timely manner or incorrect reports of possible new trackers and ads, are a root cause for lower performance. Chen et al. (Chen et al., 2021) highlighted evasion methods to circumvent blockers and filter lists: (i) changing URLs, (ii) embedding code directly into the website with the script tag, and (iii) combining functional and tracking code in one file. In summary, using filter lists for our ground truth is not without problems and directly affect the performance of our classifiers.
7.1. Limitations
Firstly, the analysis of response headers to fill the identified research gap has its strengths as well as its weaknesses. Headers represent only a part of an HTTP message, and research has shown that HTTP bodies are also relevant when it comes to understanding and identifying trackers (Li et al., 2015; Guarino. et al., 2020). From a broader point of view, only one part of application layer traffic is analyzed, disregarding other parts of web traffic, e.g. transport layer data.
Secondly, the data set could have been processed differently across the pipeline. As above-mentioned, the data sets are imbalanced, with a 72:28 non-tracker-to-tracker ratio for Chrome and Firefox data. It is not a severe form of imbalance, but it might be problematic for some classification algorithms. They could be biased against the majority class, and it might be possible to consistently predict the non-tracker label and still observe good metrics. Several resampling techniques exist, simple methods, e.g. over- and undersampling (Hernandez et al., 2013), or more complex methods, e.g. SMOTE (Chawla et al., 2002), or one of its variations (Han et al., 2005; Mukherjee and Khushi, 2021). Additional measures include using ML models, which handle imbalanced data sets well (Juez-Gil et al., 2021), or anomaly detection (Park et al., 2018). In summary, this aspect of the data pipeline has not been implemented with any of the aforementioned methods, and it might be interesting to explore how well these techniques may work for the present concept.
Thirdly, the labeling of the initial data sets, i.e. the quality of the ground truth. T.EX allows the implementation of different filter lists and offers three to date, namely EasyList, EasyPrivacy, and Disconnect.me. All base their functionality on extensive filter lists, i.e. rules and RegEx are used to block tracking-related HTTP requests. Updates are carried out regularly every four days for the first two (eas, 2023a, b), and on an irregular schedule for the latter (dis, 2023). Research has shown that these filter lists are not without errors regarding their rules (Snyder et al., 2020).
@@ 8. Conclusion
In this work, we trained Chrome-based classifiers on a subset of binary-encoded HTTP/S response headers to evaluate their effectiveness and usefulness in web tracker detection while retaining the natural imbalance of trackers. In addition, we measured the cross-browser effectiveness, showing that multiple Chrome-based classifiers achieve high accuracies for Firefox as well. However, large performance losses were observed for Brave. The longitudinal performance of the selected ML models deteriorated only by small margins, which highlights their potential due to the changing capabilities of trackers (Fouad et al., 2018; Bujlow et al., 2017; Englehardt and Narayanan, 2016). Response headers can be used to differentiate well between trackers and non-trackers and they provide additional information to the nature of tracking-related entities.
For our future research we plan to further investigate response headers, focusing on the values for feature engineering, or the use of Deep Learning models. Lastly, we will approach the deployment aspect two-fold, one from a user perspective in form of browser extension to test the models in a real-life scenario, and secondly as a fully automated system in the cloud that integrates a crawler to periodically update the models.

### v2

@@ 6.1. Intra-browser Performance
Figure 6. Radar charts plotting the performance of three selected classifiers across a subset of 13 metrics, illustrating that, while all models perform well on the in-distribution test set, the ET model exhibits more consistent performance across the various test sets, suggesting better generalizability. The models show a reduced performance on out-of-distribution test sets, indicating potential overfitting to the training data.
Table 4 shows the performance of our $\text{Chrome}_{22}$-trained classifiers and the baselines. Regarding AUPRC, the RF leads with a score of 0.98, closely followed by Extra Trees (ET) classifier and DT with scores of 0.977 and 0.965, respectively.
Except for the AdaBoost and GNB, all classifiers show a lower than baseline log-loss value, indicating a high prediction accuracy. The F1-score metric further strengthens the position of ET and RF, as they all score above 0.93, indicating that they are good at differentiating between the positive and negative classes (RQ2). The high MCC scores corroborate the other metrics, placing ET, RF, and DT as the most robust performers with accurate predictions across all categories in the confusion matrix. In contrast, baseline models like LR and AdaBoost perform comparatively lower across all metrics.
The gradient boosting models, specifically XGBoost, have fairly high scores, as reflected in an AUPRC score of 0.953 and an F1-Score of 0.883. However, their MCC scores were lower in contrast to ET, RF, and DT due to their recall scores.
Comparing the best-performing classifiers ET and RF to the baseline models shows that they perform significantly higher, except for DT, which only performed significantly worse in terms of log-loss. The t.ex-Graph classifiers display lower performance across all metrics.
@@ 6.2. Cross-browser and Longitudinal Performance
Figure 6 presents the cross-browser performance and longitudinal performance of our $\text{Chrome}_{22}$-trained classifiers. We focus on three selected classifiers: one baseline model, our best performing model, and one gradient boosting model.
@@
For ET and XGBoost, $\text{Chrome}_{22}$ shows high performance across most metrics, which is expected as the models would be most adapted to the data they were trained on. The performance on $\text{Chrome}_{23}$, $\text{Firefox}_{22}$, and $\text{Brave}_{22}$ varies, with a general trend of decreasing performance, indicating potential overfitting to the training data or a lack of generalizability to other datasets. The former should not be the case, as we observe less than a 5% performance drop between the in-distribution training and test set. Furthermore, we split the in-distribution data before the pre-processing, thus mitigating any data leakage and applied Repeated Stratified K-Fold cross-validation with five repeats and five folds. As shown in the previous sections, the datasets are different and trackers have varying structures, which is why we perform a cross-browser evaluation to test the generalizability.
@@
Overall, the results depicted in Figure 6 indicate that, while all models perform well on the dataset they were trained on, their ability to generalize to other data varies. ET exhibits the most consistent performance across different browser datasets, indicating better generalization but it had a lower log-loss score for $\text{Firefox}_{22}$ compared to the two. The figure underscores the importance of cross-browser testing to ensure that models are not only tuned to a specific dataset but also maintain their performance across diverse scenarios.
@@
The Brave results show a high accuracy metric for the overall performance, but the F1-Score and MCC are low and represent a more realistic performance, which is expected due to the considerable differences between the Chrome and Brave data.
@@
Our longitudinal performance analysis indicates a slight decline in the efficacy of the classifiers when transitioning from the $\text{Chrome}_{22}$ to the $\text{Chrome}_{23}$ dataset. We can observe a decrease in metrics such as AUC-ROC, Accuracy, and others. Notably, the MCC score between ET and XGBoost are getting similar. A trend which we can observe across most classifiers for $\text{Chrome}_{23}$ but also $\text{Firefox}_{22}$. However, the MCC scores of the gradient boosting models did not drop significantly more between $\text{Chrome}_{22}$ and $\text{Chrome}_{23}$, e.g., XGBoost went down from 0.838 to 0.823, whereas the ET suffered from 0.903 to 0.851. This suggests that while all models are affected by evolving data characteristics, ensemble methods, particularly some gradient boosting models, may offer a more stable predictive performance in longitudinal settings in contrast to the ET and RF. Nevertheless, ET and RF had better or similar performance compared to XGBoost. Applying isotonic calibration has only minor improvements compared to the uncalibrated models, as highlighted by Figure 9.
6.3. Feature Importance
As shown in Figure 7, the features Cross-Origin-Resource-Policy and Last-Modified were the important features overall, followed by Accept-Ranges and P3P. These four headers have distinct purposes in HTTP communication, except for P3P which is deprecated nowadays  [^8] https://www.w3.org/TR/P3P/. Although we plot the top ten features, the majority of them have low importance values below 0.1. DT, RF, and GBM share similar values, whereas the best-performing model, ET, has no features with a significantly higher value than the others.
Figure 7. The top ten features are presented in order of importance across different classifiers based on the in-distribution test set. The heatmap illustrates the normalized importance values for each feature and classifier. We calculated Permutation feature importance for LR and GNB but not for the others as we do not have any high cardinality features.
Iqbal et al. (Iqbal et al., 2022), who also used response header data for their classification as discussed in Section 2.1, calculated the information gain for each of their eight response-based features. This is only partially comparable to feature importance but enables two key comparisons: (i) ETag and especially P3P were important features as well, and (ii) they highlighted the importance of the Content-Length, Content-Type, and Set-Cookie headers, although they did not rank among the top 10 features in our classifiers.
6.4. HTTP Request Header Performance
Lastly, we compare the performance of our response header-based approach to classifiers that were trained on request headers. Comparing the request-based classifiers internally, we observe performance differences similar to those of their response-based counterparts. Overall, however, the request-based classifiers perform worse, as shown in Figure 8. Their AUPRC, F1-Score, and MCC values are close to 0.5 or less, which means that their prediction reliability is low.
Figure 8. Comparison between our response-based, request-based, and T.EX-Graph’s Random Forest classifier. The response-based classifier outperforms its counterparts across all evaluated metrics.
We also observe a pattern similar to the response-based classifiers regarding cross-browser and longitudinal performance. Over time, the classifiers perform slightly worse, e.g., with the MCC for the ET going down from 0.42 to 0.404. Regarding $\text{Firefox}_{22}$ data, the precision improved but the recall decreased. Nevertheless, the MCC is only 0.213, and the AUPRC is 0.458 for the ET, indicating that the classifier has no discriminative power. In the case of Brave, the classifiers are essentially ineffective and exhibit a low F1-score and MCC.
Among the most important features across the majority of classifiers are X-Client-Data, Content-Type, Origin, and Cookie. The first two are especially relevant, with values ranging from 0.2 to 0.566, excluding GNB, AdaBoost, and LightGBM. Although Content-Type is both a request and response header, it was not among the top-ten most important features for the response classifiers. The most important feature for the request-based classifiers, X-Client-Data, is used by Chrome to send an identifier to certain Alphabet-affiliated domains, e.g., doubleclick.net or googleadservices.com  [^9] https://github.com/chromium/chromium/blob/master/components/variations/net/variations_http_headers.cc, which are prevalent in our datasets. We observe similar importance values when we use an out-of-distribution dataset like $\text{Firefox}_{22}$.
@@ 7. Discussion
Tracking affects users across all three browsers but varies to a certain degree, e.g., Brave had the least amount of tracking-related HTTP messages compared to Chrome and Firefox. This can be attributed to Brave’s extensive built-in privacy protection mechanisms.
First and foremost, the application of binarization resulted in surprisingly good results, i.e., the mere presence of a header value is sufficient and extensive feature engineering is not required from a tracker detection standpoint. Our approach works well in this imbalanced setting, although it might be problematic for some classification algorithms without additional measures, such as setting class weights or rebalancing the class distribution. Additional measures include using ML models, which handle imbalanced data sets well (Juez-Gil et al., 2021), or anomaly detection (Park et al., 2018). Thus, although this has not been considered in our pipeline, it might be interesting to explore how well these measures may improve the performance of our proposed approach.
Spoofing of header values to circumvent our approach should not be a problem, as the values are not taken into account. However, this limits the models’ explainability in the sense that we lose the actual header values that might provide us with relevant information to understand trackers.
Servers can modify a response by adding or removing headers before responding to a client. The latter could impact the effectiveness of our classifiers, as they require the presence of headers. Because analyzing the characteristics and use cases of each header in our datasets is unfeasible, we focus on the top ten headers from our feature importance analysis to assess the impact of their removal on our classifiers’ ability to correctly identify trackers.
For this, we must differentiate between deprecated and actively used headers. The former may be removed by servers at any time and might prevent a positive detection by our classifiers. Examples include the Pragma header, which provides caching directives and has been superseded by the Cache-Control header, and the P3P header representing privacy practices.
Headers belonging to the latter category cannot be removed as easily. Security headers, such as X-Frame-Options, Cross-Origin-Resource-Policy, and Content-Security-Policy-Report-Only, are used to defend against various attacks like click-jacking, cross-site script inclusion attacks, or cross-site request forgery. Of equal importance are headers related to Cross-Origin-Resource-Sharing, such as Access-Control-Allow-Credentials that allows the inclusion of, e.g., credentials for HTTP authentication or cookies in cross-origin requests. Other vital headers include Last-Modified, Accept-Ranges, Etag, and Content-Encoding, which are relevant for compression, caching, or downloads. Excluding any of the aforementioned headers would have consequences leading to decreased security, performance, or functionality. Therefore, it would not be in the interest of web services to exclude these headers.
Our cross-browser approach reveals three things: firstly, the $\text{Chrome}_{22}$ dataset already captures most of the tracker structure. Secondly, due to the resulting effectiveness of a $\text{Chrome}_{22}$-trained classifier on out-of-distribution data, such as $\text{Firefox}_{22}$ and $\text{Chrome}_{23}$, it may not be necessary to re-train classifiers for these datasets. Thirdly, applying our approach to browsers like Brave that filter web tracking by default has a strong negative performance impact because the classifiers trained on unfiltered data are not able to effectively classify the modified structure of trackers in the filtered data.
The results of our clustering shown in Figure 5 reveal that trackers are more dispersed than non-trackers, which might be an argument for browser-specific models, as these might be able to better distinguish between trackers. Another challenge is the overlap between trackers and non-trackers, which further impedes correct classification. Feature engineering taking the header values into account might be a valuable approach to reduce the overlap, as header values can differ between trackers and non-trackers (see Section 4). A data-driven and large-scale in-depth analysis of header values and tracker structures across browsers is needed to explore this approach further.
Our comparison of request- and response-based classifiers revealed that request-based classifiers perform worse overall, which might be attributed to their comparatively small feature space, i.e., a smaller number of unique headers. This weaker performance suggests that request headers are less informative for the detection of web trackers than response headers. Nevertheless, the request-based classifiers correctly identify some trackers that the response-based classifiers missed, and vice-versa, as highlighted by the example of ET in Table 5. Thus, future work could explore combinations of both approaches as part of a multi-modal ensemble model.
Predicted Label by ET True Label Count Total
Response Request
NT NT T 3777 103357
NT NT NT 99580 103357
T T NT 404 7826
T T T 7422 7826
NT T T 729 19257
NT T NT 18528 19257
T NT NT 1991 41304
T NT T 39313 41304
Table 5. Comparison of the predicted and true labels for the request- and response-based ET classifier based on the in-distribution test set.
Utilizing the models for automated web tracker detection in real-life application scenarios is possible in two scenarios. First, HTTP response information is unavailable in an online (or live) detection scenario to block communication with web trackers. In this scenario, our model can enable or support reinforced learning of a complementary model on whose prediction blocking of HTTP communication is decided. This process, e.g., may imply manual inspection of HTTP communication where the predictions of our model are considered in cases where the primary model is wrong or indecisive. Second, as filter lists are still fundamental for web privacy protection, the manual process of generating filter rules can be optimized or outright replaced by utilizing our model. In this scenario, existing data sets or data from volunteers containing HTTP communication are used to identify web tracking requests. Filter rules for those can be dynamically generated and appended to conventional filter lists, which are already used.
@@
As one of the oldest methods for blocking trackers (Mazel et al., 2019), filter lists have proven to be reliable and effective. However, they have to be updated constantly, which is commonly done manually. A longitudinal study of nine years by Alrizah et al. (Alrizah et al., 2019) showed that human errors and delayed updates are a root cause for lower performance. Chen et al. (Chen et al., 2021) highlighted evasion methods to circumvent blockers and filter lists: (i) changing URLs, (ii) embedding code directly into the website with the script-tag, and (iii) combining functional and tracking code in one file. In summary, using filter lists for our ground truth is not without problems and may impact the evaluation of our classifiers. Nevertheless, they present a popular ground truth in research to label datasets for training ML models.
@@ 8. Conclusion
In this work, we train $\text{Chrome}_{22}$-based classifiers on a subset of response headers to evaluate their effectiveness and usefulness in web tracker detection while retaining the natural imbalance of trackers. Our two best classifiers RF and ET achieve scores ranging from 0.9 to 0.98. In addition, we measure the cross-browser effectiveness, showing that multiple $\text{Chrome}_{22}$-based classifiers achieve high AUPRC but moderate MCC values for $\text{Firefox}_{22}$. However, we also observe large performance losses for $\text{Brave}_{22}$.
The longitudinal performance of the selected ML models deteriorates only by small margins, which highlights their potential due to the changing capabilities of trackers (Fouad et al., 2018; Bujlow et al., 2017; Englehardt and Narayanan, 2016). Still, periodic re-training might be necessary if the performance decreases significantly.
We observe that response headers represent characteristics of web trackers which were previously under-explored by past research. They can be used to differentiate well between trackers and non-trackers and provide additional information on the nature of tracking-related entities. Our data-driven approach combined with binarization reveales new and confirms previously identified features. Applying the same methodology to request headers results in a weaker performance, thus strengthening the feasibility of using response headers for web tracker detection. Moreover, headers represent only a part of an HTTP message, and research has shown that the message bodies are also relevant when it comes to understanding and identifying trackers (Li et al., 2015; Guarino. et al., 2020). Emerging technologies like Large Language Models could be used to analyze bodies in an efficient and meaningful way.
In future research, we plan to expand on our approach of leveraging response headers for web tracker detection by considering header values in feature engineering and exploring the viability of deep-learning approaches based on response header data. Additionally, we intend to integrate our approach with existing solutions to develop a holistic web tracking detection system.
Acknowledgements.
We thank our anonymous reviewers for their feedback.